# The temporary impact of permanent hiring incentives: Evidence from Italy

### v1

@@ 6 Conclusions
Appendix A Covariate distribution
TABLE A.1: Covariate balance Male Number of renewals Potential duration Agricolture Industry Services Italian citizenship Eligible × (t = -4) 0.008-0.008 $0.001$ 0.139-0.139 $0.003$ 0.005-0.005 $0.002$ 0.032-0.032 ($0.023$) ($0.004$) ($0.151$) ($0.010$) ($0.009$) ($0.013$) ($0.021$) Eligible × (t = -3) $0.007$ $0.001$ 0.292-0.292** 0.006-0.006 0.007-0.007 $0.013$ 0.032-0.032 ($0.025$) ($0.004$) ($0.138$) ($0.011$) ($0.010$) ($0.013$) ($0.020$) Eligible × (t = -2) 0.018-0.018 $0.000$ 0.320-0.320* 0.009-0.009 0.013-0.013 $0.022$ 0.052-0.052** ($0.024$) ($0.005$) ($0.177$) ($0.013$) ($0.011$) ($0.018$) ($0.026$) Eligible × (t = 0) 0.028-0.028 0.004-0.004 0.216-0.216 $0.007$ 0.006-0.006 0.001-0.001 $0.022$ ($0.023$) ($0.006$) ($0.170$) ($0.010$) ($0.010$) ($0.013$) ($0.021$) Eligible × (t = 1) 0.037-0.037 0.009-0.009 $0.045$ $0.006$ 0.007-0.007 $0.001$ 0.040-0.040* ($0.028$) ($0.005$) ($0.144$) ($0.011$) ($0.010$) ($0.015$) ($0.021$) Standard errors in parentheses. Significance level: $^{*}p<0.1$, $^{**}p<0.05$, $^{***}p<0.01$
@@ Appendix A Covariate distribution
Appendix B Difference in Differences
TABLE B.1: Difference in differences with spillover checks Control Age 37 Control Age 38 Control Age 39 Interaction with share eligiblet-1 Eligible $\times$ (t=0) 0.0108∗∗ (0.0055) 0.0139∗∗∗ (0.0054) 0.0120∗∗ (0.0055) 0.017∗∗ (0.007) Eligible $\times$ Share eligiblet-1 $\times$ (t=0) -0.016 (0.012) Eligible $\times$ (t=-2) 0.0069 (0.0048) 0.0057 (0.0048) -0.0029 (0.0050) 0.011 (0.007) Eligible $\times$ (t=-3) -0.0015 (0.0048) 0.0036 (0.0048) 0.0033 (0.0049) 0.011 (0.007) Eligible $\times$ (t=-4) -0.0001 (0.0042) 0.0017 (0.0043) 0.0027 (0.0042) Eligible $\times$ (t=1) 0.0080 (0.0057) 0.0116∗∗ (0.0057) 0.0094 (0.0057) 0.017∗∗ (0.007) Eligible $\times$ Share eligiblet-1 $\times$ (t=-3) -0.022 (0.012) Eligible $\times$ Share eligiblet-1 $\times$ (t=-2) -0.022∗∗ (0.011) Eligible $\times$ Share eligiblet-1 $\times$ (t=1) -0.019 (0.013) t=-2 0.0008 (0.0035) 0.0024 (0.0037) 0.0113∗∗∗ (0.0040) 0.014∗∗∗ (0.005) t=-3 0.0201∗∗∗ (0.0040) 0.0154∗∗∗ (0.0037) 0.0162∗∗∗ (0.0039) 0.027∗∗∗ (0.005) t=-4 0.0000 (0.0033) -0.0013 (0.0033) -0.0019 (0.0033) t=0 0.0157∗∗∗ (0.0039) 0.0127∗∗∗ (0.0039) 0.0146∗∗∗ (0.0039) 0.017∗∗∗ (0.005) t=1 0.0335∗∗∗ (0.0046) 0.0301∗∗∗ (0.0044) 0.0323∗∗∗ (0.0045) 0.031∗∗∗ (0.005) Share eligiblet-1 $\times$ (t=-3) -0.028∗∗∗ (0.009) Share eligiblet-1 $\times$ (t=-2) -0.026∗∗∗ (0.008) Share eligiblet-1 $\times$ (t=0) -0.016∗∗ (0.008) Share eligiblet-1 $\times$ (t=1) -0.006 (0.009) Eligible $\times$ Share eligiblet-1 0.014 (0.008) Observations 72157 70475 69468 62786
@@ Appendix B Difference in Differences
Table B.1:
Appendix C Regression in Discontinuity and Difference in Discontinuity
TABLE C.1:RDD and RD-DiD regressions RDD base RDD covariates RD-DiD base RD-DiD covariates Eligible × s × (t=0) -0.0002 (0.0006) -0.0002 (0.0006) Eligible × s -0.0003 (0.0005) -0.0002 (0.0005) -0.0000 (0.0004) -0.0000 (0.0004) Eligible × s × (t=-2) 0.0000 (0.0005) 0.0000 (0.0005) Eligible × s × (t=-3) 0.0005 (0.0006) 0.0004 (0.0006) Eligible × s × (t=-4) 0.0004 (0.0005) 0.0004 (0.0005) Eligible × s × (t=1) -0.0014* (0.0008) -0.0014 (0.0008) Eligible 0.0002 (0.0003) 0.0002 (0.0003) -0.0002 (0.0002) -0.0002 (0.0002) s 0.0005 (0.0004) 0.0004 (0.0004) 0.0002 (0.0003) 0.0002 (0.0003) Eligible × (t=-2) 0.0000 (0.0002) 0.0000 (0.0002) Eligible × (t=-3) -0.0002 (0.0004) -0.0002 (0.0004) Eligible × (t=-4) 0.0001 (0.0002) 0.0001 (0.0002) Eligible × (t=0) 0.0004 (0.0003) 0.0004 (0.0003) Eligible × (t=1) 0.0014** (0.0006) 0.0014** (0.0006) s × (t=-2) -0.0001 (0.0004) -0.0001 (0.0004) s × (t=-3) -0.0004 (0.0005) -0.0004 (0.0005) s × (t=-4) -0.0003 (0.0003) -0.0003 (0.0003) s × (t=0) 0.0002 (0.0005) 0.0002 (0.0005) s × (t=1) 0.0003 (0.0005) 0.0003 (0.0005) Observations 64432 64432 380629 380629 Covariates
@@ Appendix C Regression in Discontinuity and Difference in Discontinuity
Table C.1:
Figure C.1: Density test for the RD-DiD
@@
Appendix D Medium run effects
TABLE D.1: DiD medium-run outcomes two years after Same contract Permanent contract Same firm Same sector Days worked Eligible $\times$ (t=0) -0.004 (0.009) 0.012 (0.015) 0.013 (0.014) 0.014 (0.014) 0.012 (0.012) Eligible $\times$ (t=-4) -0.004 (0.007) -0.010 (0.015) -0.010 (0.015) -0.005 (0.014) -0.001 (0.012) Eligible $\times$ (t=-3) -0.008 (0.008) -0.010 (0.015) -0.004 (0.015) -0.006 (0.015) -0.002 (0.013) Eligible $\times$ (t=-2) -0.010 (0.008) -0.028* (0.015) -0.027* (0.015) -0.030** (0.014) -0.018 (0.012) t=-4 -0.043*** (0.006) -0.010 (0.012) -0.009 (0.011) -0.019* (0.011) -0.009 (0.009) t=-3 -0.029*** (0.006) -0.001 (0.011) -0.008 (0.011) -0.019* (0.011) -0.021** (0.009) t=-2 -0.019** (0.006) -0.030*** (0.011) -0.033*** (0.011) -0.043*** (0.011) -0.040*** (0.009) t=0 0.012 (0.006) -0.007 (0.011) -0.005 (0.010) 0.001 (0.010) 0.014* (0.008) Eligible 0.013* (0.006) 0.020* (0.011) 0.018* (0.010) 0.021** (0.010) 0.014 (0.009) Num. Obs. 63646 35683 35683 35683 35683
@@ Appendix D Medium run effects
Table D.1:
TABLE D.2: DiD medium-run outcomes three years after Same contract Permanent contract Same firm Same sector Days worked Eligible $\times$ (t=0) 0.002 (0.008) 0.013 (0.015) 0.014 (0.014) 0.015 (0.014) 0.007 (0.011) Eligible $\times$ (t=-4) -0.001 (0.006) 0.000 (0.015) 0.000 (0.015) 0.004 (0.014) 0.008 (0.011) Eligible $\times$ (t=-3) -0.000 (0.007) 0.004 (0.016) 0.008 (0.015) 0.006 (0.015) 0.003 (0.012) Eligible $\times$ (t=-2) -0.004 (0.007) -0.024 (0.015) -0.023 (0.015) -0.029* (0.015) -0.014 (0.012) t=-4 -0.036*** (0.005) -0.022* (0.012) -0.026** (0.011) -0.044*** (0.011) -0.045*** (0.009) t=-3 -0.026*** (0.005) -0.003 (0.012) -0.014 (0.011) -0.031*** (0.011) -0.038*** (0.009) t=-2 -0.019*** (0.005) -0.004 (0.011) -0.007 (0.011) -0.022** (0.011) -0.033*** (0.009) t=0 0.012* (0.005) 0.003 (0.011) -0.003 (0.011) -0.003 (0.010) -0.004 (0.008) Eligible 0.008 (0.005) 0.016 (0.011) 0.016 (0.010) 0.018* (0.010) 0.013 (0.008) Num. Obs. 63646 35683 35683 35683 35683
@@
Table D.2:

### v2

@@ Acknowledgements
Appendix A Concurrent policies
The Decreto Dignità (Decree-Law No. 87 of 12 July 2018) was adopted by the Italian government and subsequently converted into law (Law No. 96 of 9 August 2018) after parliamentary scrutiny and amendments. It was initially approved in mid-July and became effective shortly thereafter. At this stage, the new rules on fixed-term employment contracts applied primarily to newly stipulated contracts. Subsequently, through conversion into law (Law 9 August 2018, n. 96) and accompanying ministerial/circular clarifications, it was established that, from 31 October 2018 onward, the reform’s provisions would extend their reach to also cover renewals or extensions of contracts already in place as of 14 July 2018. Thus, from November 1, 2018, the new discipline was applied comprehensively.
The measure aimed to reduce job precariousness by tightening rules on fixed-term contracts and strengthening protections for permanent employment. Among the major modifications introduced: (i) the maximum duration of fixed-term contracts was reduced from 36 to 24 months, (ii) the maximum number of fixed-term contract renewals was cut from five to four, (iii) the requirement to report an explicit causal reason (“causale”) for the termination of a contract was extended to fixed-term contracts exceeding 12 months; (iv) social contributions labour costs were increased from 1,4% to 1,9%.
Several caveats apply, limiting the impact of this policy on our estimates. Firstly, it is essential to note that the new Decree-Law does not directly impact conversion per se and does not affect the 35-year-old age threshold that defines the group of eligible beneficiaries. The introduction of this policy can, then, at worst, moderate the effects of the incentives by increasing overall conversions. Secondly, the policy only became effective in late 2018, meaning that only 36% of contracts started after the initial partial implementation of the decree.
Figure A.1: Conversions by months from 2017 to 2019
The plot displays the monthly conversions divided by people aged 33 and 35. The vertical solid lines indicate the boundaries between years, while the two dotted ones mark the partial (first) and full (second) implementation of the Decreto Dignità.
In light of these considerations, we can compare contract conversions before and after the introduction of the concurrent policy. Figure A.1 shows that seasonal factors predominantly explain the observed dynamics, as the three years follow a comparable monthly pattern, though with a positive underlying trend. The main differences in conversion rates between the two age groups emerge in the early months of the year 2018, before the approval of the Degree. This pattern suggests that, once seasonality is properly controlled for, the short-term impact of the Decreto Dignità on the likelihood of contract conversion appears to be limited. Regression results also seem to point in the same direction. In Table A.1 we report the estimates of the treatment interacted with a dummy variable equal to one for individuals born before August 1st. This variable allows us to examine potential heterogeneity between those who were simultaneously eligible for both reforms, the incentives and the Decreto Dignità, and other individuals. We report only the relevant coefficients, which indicate no meaningful difference between the two subgroups, suggesting a negligible effect on conversions of the more recent Decree.
Nonetheless, we cannot entirely rule out the possibility that, in the medium term, the Decreto Dignità may have either moderated or mitigated the observed incentive effect. The absence of interaction effects in the short term suggests that the results should also hold in the medium term. Hence, while some degree of moderation cannot be excluded, the identification strategy seems to remain valid, as does the internal validity of the analysis.
TABLE A.1: Interaction with Decreto Dignità Covariate Model Eligible 0.005-0.005 8 ($0.005$) Born before August 0.006-0.006 7 ($0.004$) Eligible × (t = 0) 0.0130.013* 2 ($0.007$) Eligible × Born before August × (t = 0) $0.002$ ($0.010$) Num.Obs. 113 576113\,576 Std.Errors by: Firm Covariates Standard errors are in parentheses. Significance levels: * p << 0.1, ** p << 0.05, *** p << 0.01
@@ Appendix A Concurrent policies
Appendix B Covariates distribution
TABLE B.1: Covariates balance Male Number of renewals Potential duration Agricolture Industry Services Italian citizenship Eligible × (t = -4) 0.008-0.008 $0.001$ 0.139-0.139 $0.003$ 0.005-0.005 $0.002$ 0.032-0.032 ($0.023$) ($0.004$) ($0.151$) ($0.010$) ($0.009$) ($0.013$) ($0.021$) Eligible × (t = -3) $0.007$ $0.001$ 0.292-0.292** 0.006-0.006 0.007-0.007 $0.013$ 0.032-0.032 ($0.025$) ($0.004$) ($0.138$) ($0.011$) ($0.010$) ($0.013$) ($0.020$) Eligible × (t = -2) 0.018-0.018 $0.000$ 0.320-0.320* 0.009-0.009 0.013-0.013 $0.022$ 0.052-0.052** ($0.024$) ($0.005$) ($0.177$) ($0.013$) ($0.011$) ($0.018$) ($0.026$) Eligible × (t = 0) 0.028-0.028 0.004-0.004 0.216-0.216 $0.007$ 0.006-0.006 0.001-0.001 $0.022$ ($0.023$) ($0.006$) ($0.170$) ($0.010$) ($0.010$) ($0.013$) ($0.021$) Eligible × (t = 1) 0.037-0.037 0.009-0.009 $0.045$ $0.006$ 0.007-0.007 $0.001$ 0.040-0.040* ($0.028$) ($0.005$) ($0.144$) ($0.011$) ($0.010$) ($0.015$) ($0.021$) Standard errors in parentheses. Significance level: $^{*}p<0.1$, $^{**}p<0.05$, $^{***}p<0.01$
Table B.1:
Appendix C Difference in Differences
TABLE C.1: Difference in differences with spillover checks Control Age 37 Control Age 38 Control Age 39 Interaction with share eligiblet-1 Eligible $\times$ (t=0) 0.0108∗∗ (0.0055) 0.0139∗∗∗ (0.0054) 0.0120∗∗ (0.0055) 0.017∗∗ (0.007) Eligible $\times$ Share eligiblet-1 $\times$ (t=0) -0.016 (0.012) Eligible $\times$ (t=-2) 0.0069 (0.0048) 0.0057 (0.0048) -0.0029 (0.0050) 0.011 (0.007) Eligible $\times$ (t=-3) -0.0015 (0.0048) 0.0036 (0.0048) 0.0033 (0.0049) 0.011 (0.007) Eligible $\times$ (t=-4) -0.0001 (0.0042) 0.0017 (0.0043) 0.0027 (0.0042) Eligible $\times$ (t=1) 0.0080 (0.0057) 0.0116∗∗ (0.0057) 0.0094 (0.0057) 0.017∗∗ (0.007) Eligible $\times$ Share eligiblet-1 $\times$ (t=-3) -0.022 (0.012) Eligible $\times$ Share eligiblet-1 $\times$ (t=-2) -0.022∗∗ (0.011) Eligible $\times$ Share eligiblet-1 $\times$ (t=1) -0.019 (0.013) t=-2 0.0008 (0.0035) 0.0024 (0.0037) 0.0113∗∗∗ (0.0040) 0.014∗∗∗ (0.005) t=-3 0.0201∗∗∗ (0.0040) 0.0154∗∗∗ (0.0037) 0.0162∗∗∗ (0.0039) 0.027∗∗∗ (0.005) t=-4 0.0000 (0.0033) -0.0013 (0.0033) -0.0019 (0.0033) t=0 0.0157∗∗∗ (0.0039) 0.0127∗∗∗ (0.0039) 0.0146∗∗∗ (0.0039) 0.017∗∗∗ (0.005) t=1 0.0335∗∗∗ (0.0046) 0.0301∗∗∗ (0.0044) 0.0323∗∗∗ (0.0045) 0.031∗∗∗ (0.005) Share eligiblet-1 $\times$ (t=-3) -0.028∗∗∗ (0.009) Share eligiblet-1 $\times$ (t=-2) -0.026∗∗∗ (0.008) Share eligiblet-1 $\times$ (t=0) -0.016∗∗ (0.008) Share eligiblet-1 $\times$ (t=1) -0.006 (0.009) Eligible $\times$ Share eligiblet-1 0.014 (0.008) Observations 72157 70475 69468 62786
@@ Appendix C Difference in Differences
Table C.1:
Appendix D Regression in Discontinuity and Difference in Discontinuity
TABLE D.1: RDD and RD-DiD regressions RDD base RDD covariates RD-DiD base RD-DiD covariates Eligible × Share eligible -0.0003 (0.0005) -0.0002 (0.0005) -0.0000 (0.0004) -0.0000 (0.0004) Eligible 0.0002 (0.0003) 0.0002 (0.0003) -0.0002 (0.0002) -0.0002 (0.0002) Share eligible 0.0005 (0.0004) 0.0004 (0.0004) 0.0002 (0.0003) 0.0002 (0.0003) Eligible × Share eligible × (t=0) -0.0002 (0.0006) -0.0002 (0.0006) Eligible × Share eligible × (t=-2) 0.0000 (0.0005) 0.0000 (0.0005) Eligible × Share eligible × (t=-3) 0.0005 (0.0006) 0.0004 (0.0006) Eligible × Share eligible × (t=-4) 0.0004 (0.0005) 0.0004 (0.0005) Eligible × Share eligible × (t=1) -0.0014* (0.0008) -0.0014 (0.0008) Eligible × (t=-2) 0.0000 (0.0002) 0.0000 (0.0002) Eligible × (t=-3) -0.0002 (0.0004) -0.0002 (0.0004) Eligible × (t=-4) 0.0001 (0.0002) 0.0001 (0.0002) Eligible × (t=0) 0.0004 (0.0003) 0.0004 (0.0003) Eligible × (t=1) 0.0014** (0.0006) 0.0014** (0.0006) Share eligible × (t=-2) -0.0001 (0.0004) -0.0001 (0.0004) Share eligible × (t=-3) -0.0004 (0.0005) -0.0004 (0.0005) Share eligible × (t=-4) -0.0003 (0.0003) -0.0003 (0.0003) Share eligible × (t=0) 0.0002 (0.0005) 0.0002 (0.0005) Share eligible × (t=1) 0.0003 (0.0005) 0.0003 (0.0005) Observations 64432 64432 380629 380629 Covariates
@@ Appendix D Regression in Discontinuity and Difference in Discontinuity
Table D.1:
Figure D.1: Density test for the RD-DiD
@@
Appendix E Medium run effects
TABLE E.1: DiD medium-run outcomes two years after Same contract Permanent contract Same firm Same sector Days worked Eligible $\times$ (t=0) -0.004 (0.009) 0.012 (0.015) 0.013 (0.014) 0.014 (0.014) 0.012 (0.012) Eligible $\times$ (t=-4) -0.004 (0.007) -0.010 (0.015) -0.010 (0.015) -0.005 (0.014) -0.001 (0.012) Eligible $\times$ (t=-3) -0.008 (0.008) -0.010 (0.015) -0.004 (0.015) -0.006 (0.015) -0.002 (0.013) Eligible $\times$ (t=-2) -0.010 (0.008) -0.028* (0.015) -0.027* (0.015) -0.030** (0.014) -0.018 (0.012) t=-4 -0.043*** (0.006) -0.010 (0.012) -0.009 (0.011) -0.019* (0.011) -0.009 (0.009) t=-3 -0.029*** (0.006) -0.001 (0.011) -0.008 (0.011) -0.019* (0.011) -0.021** (0.009) t=-2 -0.019** (0.006) -0.030*** (0.011) -0.033*** (0.011) -0.043*** (0.011) -0.040*** (0.009) t=0 0.012 (0.006) -0.007 (0.011) -0.005 (0.010) 0.001 (0.010) 0.014* (0.008) Eligible 0.013* (0.006) 0.020* (0.011) 0.018* (0.010) 0.021** (0.010) 0.014 (0.009) Num. Obs. 63646 35683 35683 35683 35683
@@ Appendix E Medium run effects
Table E.1:
TABLE E.2: DiD medium-run outcomes three years after Same contract Permanent contract Same firm Same sector Days worked Eligible $\times$ (t=0) 0.002 (0.008) 0.013 (0.015) 0.014 (0.014) 0.015 (0.014) 0.007 (0.011) Eligible $\times$ (t=-4) -0.001 (0.006) 0.000 (0.015) 0.000 (0.015) 0.004 (0.014) 0.008 (0.011) Eligible $\times$ (t=-3) -0.000 (0.007) 0.004 (0.016) 0.008 (0.015) 0.006 (0.015) 0.003 (0.012) Eligible $\times$ (t=-2) -0.004 (0.007) -0.024 (0.015) -0.023 (0.015) -0.029* (0.015) -0.014 (0.012) t=-4 -0.036*** (0.005) -0.022* (0.012) -0.026** (0.011) -0.044*** (0.011) -0.045*** (0.009) t=-3 -0.026*** (0.005) -0.003 (0.012) -0.014 (0.011) -0.031*** (0.011) -0.038*** (0.009) t=-2 -0.019*** (0.005) -0.004 (0.011) -0.007 (0.011) -0.022** (0.011) -0.033*** (0.009) t=0 0.012* (0.005) 0.003 (0.011) -0.003 (0.011) -0.003 (0.010) -0.004 (0.008) Eligible 0.008 (0.005) 0.016 (0.011) 0.016 (0.010) 0.018* (0.010) 0.013 (0.008) Num. Obs. 63646 35683 35683 35683 35683
@@
Table E.2:
Appendix F Heterogeneity
F.1 Short run
Figure F.1:
Panel a - Macrosector
Panel b - Skill Level
Panel c - Firm Size
Panel d - Potential Duration
Panel e - Gender
F.2 Medium run - 1 year forward
Figure F.2:
Panel a - Macrosector
Panel b - Skill Level
Panel c - Firm Size
Panel d - Potential Duration
Panel e - Gender
F.3 Medium run - 4 year forward
Figure F.3:
Panel a - Macrosector
Panel b - Skill Level
Panel c - Firm Size
Panel d - Potential Duration
Panel e - Gender